Solve the following compound inequality and show the solution on a number line and in interval notation | x-5 | - 4=0
Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

|x-5| - 4=0
|x-5| = 4
-----
x - 5 = 4
x = 9
------------
-(x-5) = 4
x-5 = -4
x = 1

Answer by ikleyn(52852)

.
There is NO inequality in this post.